OpenStudy (anonymous):

@jigglypuff314 can you help me

jigglypuff314 (jigglypuff314):

I'll see if I can :) what's your question?

OpenStudy (anonymous):

Thank you .. How can i simplify 176 over 14.?

OpenStudy (anonymous):

And can you draw it out for me

jigglypuff314 (jigglypuff314):

well both 176 and 14 can be divided by 2 right? so 176/2 = 88 and 14/2 = 7 therefore fraction 176 over 14 = 88 over 7

OpenStudy (anonymous):

But I have to Reduce it to its lowest fraction

jigglypuff314 (jigglypuff314):

I think that is the lowest... http://www.wolframalpha.com/input/?i=176%2F14

OpenStudy (whpalmer4):

If you are reducing a fraction such as \[\frac{176}{14}\]and you get to the point where either the numerator or the denominator is a prime number (such as 7), which has no factors besides 1 and itself, you are done. You are also done if the numerator and denominator have no factors in common, even though they may have factors. For example, \[\frac{25}{9} = \frac{5*5}{3*3}\]Both 25 and 9 have factors, but they have no factors in common, so you cannot reduce the fraction.